# Geometric, topological and dynamical properties of conformally symplectic systems, normally hyperbolic invariant manifolds, and scattering maps

### v1

@@ 1.1. Overview
Conformally symplectic maps and their invariant manifolds present intrinsic mathematical interest, e.g., [AA24, Ban02, Lee43, Vai76, Vai85]. They also arise in applications, such as mechanical systems with friction proportional to velocity, e.g., [CCdlL22, MOR14, BC09], as Euler-Lagrange equations of exponentially discounted systems – which appear in financial models and control theory – e.g., [Ben88], or thermostat systems, e.g. [WL98]
@@
The case $\eta=1$ corresponds to symplectic maps. The case $\eta\neq 1$ has some similarities with the symplectic case but also important differences111 The limit $\eta\to 1$ is a singular limit, as the properties of the system change dramatically.. Notably, the symplectic volume changes under iteration by a factor, $f^{*}\left(\omega^{\wedge d/2}\right)=\eta^{d/2}\left(\omega^{\wedge d/2}\right)$. So, the only invariant objects are of volume 0 or $\infty$. The volume 0 case includes interesting objects such as Birkhoff attractors [AHV24], however, in this paper we will concentrate on invariant manifolds of infinite volume. In the infinite-volume setting, uniform boundedness properties of differentiable objects are not straightforward, and they must be carefully formulated and handled.
@@
On the other hand, we do not need that the non-degeneracy properties of $\omega$ are uniform (that is, we can allow that $|\omega^{\wedge d/2}|$ is not equivalent to the Riemannian volume). Many of the results of the main Theorems 3.1 and 3.3 (not the pairing rules obtained in Theorem 3.1) work even when $\omega$ is a presymplectic form (see Theorem 3.9).
Many of the remarkable properties of symplectic dynamics extend, often with suitable modifications, to the conformally symplectic setting. Examples include KAM [CCdlL13, CCdlL22], and, under convexity assumptions, Aubry-Mather theory [MS17] and Hamilton-Jacobi theory [Gom08]. On the other hand, new phenomena such as attractors appear for conformally symplectic systems. In conformally symplectic (but not symplectic) systems it is impossible to have invariant manifolds finite non-zero volume.
@@ 1.2.1. Topology and conformally symplectic dynamics
    Notice that, even if we assume that $\omega$ is bounded, it is not natural – and often impossible – to assume $\alpha$ is bounded. See Section 9.4. In the conformally symplectic case, it happens often that a map is exact for some action forms but not for others (see Section 5.1). This has implication on the de Rham cohomology of the manifold. For two action forms $\alpha,\tilde{\alpha}$ for $\omega$, we have $d(\alpha-\tilde{\alpha})=0$ but not necessarily $\tilde{\alpha}-\alpha=dG$. For any action form $\alpha$, $\alpha+dG$ is an action form too, and if $f$ is exact for $\alpha$, then $f$ is also exact for $\tilde{\alpha}$. We call the addition of $dG$ to an action form a gauge transformation. We study systematically the effect of gauge transformations on primitive functions.
  • In some manifolds, there are relations between the algebraic topology and the conformal factors $\eta$. Consequently, on these manifolds, the possible values of $\eta$ for non-exact $\omega$ are algebraic numbers. See Section 4.2.
    This solves a question raised in [AF24].
@@ 1.2.2. Vanishing lemmas
    $|\omega(x)(u,v)|\leq\eta^{-n}\|Df^{n}(x)u\|\|Df^{n}(x)v\|.$
    From this, we can derive vanishing lemmas stating that the form $\omega$ vanishes on subbundles of tangent vectors whose growth rates satisfy certain relations with respect to $\eta$.
@@
    These allows us to prove global versions for NHIM of the widely studied pairing rules [Des88, DM96, WL98] for periodic orbits of Lyapunov exponents.
@@ 1.2.3. NHIMs for conformally symplectic systems
We recall that a NHIM is an invariant manifold such that there are gaps between the rates of growth of vectors tangent to the manifold and the rates of growth in the stable and unstable bundles that span the normal bundle. See [Fen71, Fen74, Fen77, Pes04, BLZ08], and also Appendix A.
@@
It should be noted that the unboundedness of manifolds is not merely a technical inconvenience; it can give rise to new geometric phenomena, such as an NHIM folding into itself. See [Eld12, Example 3.8], reproduced here in Figure 2. To avoid such pathological examples, we have introduced an explicit assumption (U2) that the NHIM has a uniform tubular neighborhood.
@@ 1.2.4. Scattering maps for conformally symplectic systems
An important tool in the study of homoclinic excursions to a NHIM is the scattering map, introduced in [DdlLS00].
We consider a class of homoclinic excursions given by a transversal intersection of stable/unstable manifolds (the precise conditions are given in Definition 2.23 in Section 2.10). Given a homoclinic excursion $\{f^{n}(x)\}_{n\in\mathbb{Z}}$ to a NHIM $\Lambda$,
there are unique points $x_{+},x_{-}\in\Lambda$ such that
@@
where $\lambda$ is strictly bigger than the rates in the tangent space.
@@
The scattering map is a very useful tool to understand long range motions and instability. Most of the applications of the scattering map in the instability problem have been for symplectic dynamics, since it gives a global way to connect invariant objects such as whiskered tori of (possibly) different topology or dimension [DdlLS06, DdlLS16]. There are also results using the scattering map in systems with dissipation [Gra17, GdlLM22, AGMS23].
The scattering map is not itself part of the dynamics; rather, it should be regarded as a comparison between the dynamics restricted to the NHIM and the dynamics along the homoclinic excursion. Nevertheless, a surprising result [GLLMS20, GdlLS20] shows that iterations of the scattering map interspersed with iterations of the dynamics restricted to the NHIM corresponds to true orbits of the map. The basic results of the previous two papers apply to general systems, including conformally symplectic systems. See also [DGR12].
Geometric properties of the scattering map in the symplectic case were systematically studied in [DdlLS08]. In this paper we develop similar results for conformally symplectic systems, but we encounter significant differences.
@@
    The formulas consist of a finite series plus an integral term as the remainder. Similar formulas were known for symplectic twist maps, and used for numerical computations [Tab95].
    The theory behind these formulas for conformally symplectic maps has some surprises. As we pointed out, there are different action forms corresponding to gauge transformations. The gauges chosen can affect the convergence of the series.
@@
  • We provide a variational interpretation of the primitive function of the scattering map. We remark that, under certain convexity assumptions and provided the Legendre transform can be implemented, the primitive function of the scattering map coincides with the renormalized variational principle used to construct connecting orbits. [Rab08, Bes96].
@@ 2.1.1. Riemannian manifolds, differentiable maps and forms
The sets of $\mathcal{C}^{r}$ vector fields, forms, and other objects with a linear structure, have a $\mathcal{C}^{r}$ norm given by the supremum of the size of the objects and their derivatives, which makes them into a Banach space. Under suitable uniformity assumptions (see (U1)), the set of $\mathcal{C}^{r}$ diffeomorphisms is a Banach manifold [Ban97].
@@ 2.1.2. Distances and regularity of manifolds
The following definitions are standard, but we record them here since there are subtle. See also [PSW97].
@@
In this paper, we will need to study some foliations of a manifold by submanifolds. Foliations have regularity characterized by two numbers. One of them is the regularity of the leaves and another one is the dependence of the leaves on the base point along a transversal. We use the definition of [HPS77, PSW97].
@@ 2.1.3. Exponentials and connectors
It is well known [BG05, Prop 4.5.2] that on a ball $B_{\rho_{x}}(0)\subseteq T_{x}M$ of radius $0<\rho_{x}\ll 1$, the exponential mapping is a diffeomorphism and induces a system of smooth coordinates on the neighborhood $\exp_{x}(B_{\rho_{x}}(0))$ of $x$ in $M$, referred to as geodesic coordinates. For any $x\in M$, the supremum of all such radii $\rho_{x}$ is called the injectivity radius.
@@
If $x,y$ are sufficiently close (depending only on derivatives of the metric) then $S_{x}^{y}$ is invertible, and we think of it as providing an identification of the two tangent spaces. The maps $S_{x}^{y}$ are called connectors in [HPPS70], where they are used in applications to hyperbolic systems.
@@ 2.2. A standing assumption on the manifold
Hypothesis (U1’) is sufficient but not necessary to obtain (U1). To obtain that the geodesic flow (and the exponential mapping) are $\mathcal{C}^{r}$, one does not need to assume that all the derivatives of $g$ are bounded. Only that the derivatives of the Riemann tensor are bounded. This is called bounded geometry [CGT82, Eld12].
@@ A $2$-form on a $\mathcal{C}^{r}$-manifold $M$ of even dimension $d$
The form $\omega$ is exact if $\omega=d\alpha$ for some $1$-form [^2] We refer to $\alpha$ as an action form for $\omega$ following [Har00], but other names are used in the literature such as symplectic potential, Liouville form, or primitive form. $\alpha$ (not necessarily unique).
@@
A submanifold $L\subset M$ is said to be isotropic if $\omega_{\mid L}=0$, that is, for each $y\in L$, $T_{y}L\subseteq T_{y}L^{\omega}$, where $T_{y}L^{\omega}=\{v\in T_{y}M\,|\,\omega(y)(v,u)=0,\,\forall u\in T_{y}L\}$ is the symplectic orthogonal of $T_{y}L$. A submanifold $L\subset M$ is said to be coisotropic if for each $y\in L$, $T_{y}L^{\omega}\subseteq T_{y}L$. A submanifold $L\subset M$ is Lagrangian if it is both isotropic and coisotropic, or, equivalently, $\omega_{\mid L}=0$ and $\textrm{dim}(L)=\textrm{dim}(M)/2$. (See, e.g., [Wei71, Wei73].)
We will also pay attention to forms that are degenerate and have constant rank (dimension of the kernel) [Sou97].
@@
In some works [Sou97, LM87], the definition of presymplectic forms also includes the constant rank condition that $\dim K_{x}(\omega)=$ const. on open sets. In general, the constant rank is not an open condition, since a small perturbation of the form may decrease the dimension of the kernel.
@@ A $\mathcal{C}^{r}$-map $f:M\to M$ is conformally symplectic if
It is well known([LM87, Ban02]) that if $\omega$ is a two form of rank greater or equal than $4$, then, if, for a $1$-form $\beta$, we have $\beta\wedge\omega=0$, we conclude $\beta=0$.
@@
Taking $d$, $0=f^{*}(d\omega)=d(f^{*}\omega)=d\eta\wedge\omega$. Hence $d\eta=0$.
When $\eta=1$ a system satisfying (2.3) is symplectic, so the results in this paper (that do not include explicitly $\eta\neq 1$) imply the results in [DdlLS08]. When $\eta\neq 1$, the symplectic volume contracts or expands, and, therefore, invariant manifolds have volume zero or infinite. The interesting case is when the symplectic volume is infinite. In many mechanical applications, the physical friction satisfies $\eta<1$.
@@ 2.5. Exact conformally symplectic maps
In many cases of interest, there are topological reasons why the symplectic form $\omega$ has to be exact; see [AF24] and Section 4.1.
@@ 2.5.1. Gauge transformations
In general, if a map is exact for $\alpha$, it will also be exact under the gauge transformation of the action form  [^3] In electromagnetism, where the electromagnetic field is an exact 2-form exterior derivative of an electromagnetic 1-form (called vector potential), one often uses adding gradient to the vector potential [Thi97, Zan13]. The term “gauge” was introduced in [Wey51] for electromagnetic vector potentials
@@
More generally, given any construction, it is natural to ask how does it behave under gauge transformations.
@@ 2.7. Normally hyperbolic invariant manifolds
Definition 2.16 says that in a small neighborhood of $\Lambda\subset M$ of any point $x\in\Lambda$, the manifold $M$ is the product of the manifold $\Lambda$ and of the fibers of the bundles $E^{s}$ and $E^{u}$. For compact manifolds, this gives a tubular neighborhood of $\Lambda$. See [HPS77].
For unbounded manifolds, it can happen that the manifold $\Lambda$ folds back and comes close to itself. A concrete example appears in [Eld12, Example 3.8], which we illustrate in Fig. 2. The content of assumption (U2) is that this folding back of the manifold $\Lambda$ does not happen.
Figure 2. A NHIM without a uniform tubular neighborhood. After [Eld12, Example 3.8]
@@
$W^{u,s,\textrm{loc}}_{\Lambda}=\bigcup_{x\in\Lambda}W^{s,u,\textrm{loc}}_{x}$
The union in the right-hand side above is a disjoint union. So that the decomposition of $W^{u,s,\textrm{loc}}_{\Lambda}$ into strong stable leaves is a foliation.
These manifolds and their regularity are described in Appendix A (see Theorem A.1). We will take advantage of the remarkable fact that $W^{u,s,\textrm{loc}}_{\Lambda}$ admit topological characterizations, but has consequences for rates of convergence and regularity.
@@
The assumption (N) is to avoid complicated statements. If $\mu_{+}<1$, there is a unique fixed point $p\in\Lambda$, and $\Lambda$ is a submanifold of the stable manifold of pp. Many of the regularity statements obtained in the general theory of NHIMs remain true, but they are far from optimal. Also, in some estimates (derived from [DdlLS08, Proposition 15] or similar) we use bounds like $C_{1}\mu_{+}^{n}+C_{2}\mu_{+}^{2n}\leq C\mu_{+}^{2n}$, for $n\geq 0$. If we do not assume (N), different algebraic expressions for the bounds would be required for $\mu_{+}\geq 1$ and for $\mu_{+}<1$, and therefore many statements would need to distinguish between the two different cases.
@@
Remark 2.18.
Since we consider invariant manifolds for conformally symplectic systems that are unbounded (see Remark 2.12) we need to make explicitly the uniformity assumptions (U1) (or (U1’)) on the manifold $M$. Some aspects of the analysis require that the invariant manifold $\Lambda$ has a uniform tubular neighborhood. See assumption $\bf(U2)$}. For the analysis on the NHIM $\Lambda$ and its homoclinic excursions, it would suffice to study a neighborhood of $\Lambda$ and its iterates. We make assumptions on the regularity of the map $f$ in such a neighborhood. See assumptions (U3) and (U4).

### v2

@@ 1.1. Overview
Conformally symplectic maps and their invariant manifolds present intrinsic mathematical interest, e.g., [AllaisA24, Banyaga02, Lee46, Vaisman75, Vaisman85]. They also arise in applications, such as mechanical systems with friction proportional to velocity, e.g., [CallejaCL22, MargheriOR14, BiascoCh09], as Euler-Lagrange equations of exponentially discounted systems – which appear in financial models and control theory – e.g., [Bensoussan88], or thermostat systems, e.g. [wojtkowski1998conformally]
@@
The case $\eta=1$ corresponds to symplectic maps. The case $\eta\neq 1$ has some similarities with the symplectic case but also important differences [^1] The limit $\eta\to 1$ is a singular limit, as the properties of the system change dramatically.. Notably, the symplectic volume changes under iteration by a factor, $f^{*}\left(\omega^{d/2}\right)=\eta^{d/2}\left(\omega^{d/2}\right)$. So, the only invariant objects are of volume 0 or $\infty$. The volume 0 case includes interesting objects such as Birkhoff attractors [arnaud2024higherdimensionalbirkhoffattractors], however, in this paper we will concentrate on invariant manifolds of infinite volume. In the infinite-volume setting, uniform boundedness properties of differentiable objects are not straightforward, and they must be carefully formulated and handled.
@@
On the other hand, we do not need that the non-degeneracy properties of $\omega$ are uniform (that is, we can allow that $|\omega^{d/2}|$ is not equivalent to the Riemannian volume). Many of the results of the main Theorems 3.1 and 3.3 (not the pairing rules obtained in Theorem 3.1) work even when $\omega$ is a presymplectic form (see Theorem 3.9).
Many of the remarkable properties of symplectic dynamics extend, often with suitable modifications, to the conformally symplectic setting. Examples include KAM [CallejaCL13, CallejaCL22], and, under convexity assumptions, Aubry-Mather theory [maro2017aubry] and Hamilton-Jacobi theory [Gomes08]. On the other hand, new phenomena such as attractors appear for conformally symplectic systems. In conformally symplectic (but not symplectic) systems it is impossible to have invariant manifolds finite non-zero volume.
@@ 1.2.1. Topology and conformally symplectic dynamics
    Notice that, even if we assume that $\omega$ is bounded, it is not natural – and often impossible – to assume $\alpha$ is bounded. See Section 9.4. In the conformally symplectic case, it happens often that a map is exact for some action forms but not for others (see Section 5.1).
    This has relations with the de Rham cohomology of the manifold. For two action forms $\alpha,\tilde{\alpha}$ for $\omega$, we have $d(\alpha-\tilde{\alpha})=0$ but not necessarily $\tilde{\alpha}-\alpha=dG$. For any action form $\alpha$, $\alpha+dG$ is an action form too, and if $f$ is exact for $\alpha$, then $f$ is also exact for $\tilde{\alpha}$. We call the addition of $dG$ to an action form a gauge transformation. We study systematically the effect of gauge transformations on primitive functions.
  • In some manifolds, there are relations between the algebraic topology and the conformal factors $\eta$. Consequently, on these manifolds, the possible values of $\eta$ for non-exact $\omega$ are algebraic numbers. See Section 4.2. Using these relations, in Section 4.2, we present an answer to a question raised in [ArnaudF24].
@@ 1.2.2. Vanishing lemmas
    $|\omega(x)(u,v)|\leq\eta^{-n}\|\omega\|\|Df^{n}(x)u\|\|Df^{n}(x)v\|.$
    From this, we can derive vanishing lemmas stating that $\omega$ vanishes on subbundles of tangent vectors whose growth rates satisfy certain relations with respect to $\eta$.
@@
    These allows us to prove global versions for NHIM of the widely studied pairing rules [Dessler88, dettmann1996proof, wojtkowski1998conformally] for periodic orbits of Lyapunov exponents.
@@ 1.2.3. NHIMs for conformally symplectic systems
We recall that a NHIM is an invariant manifold such that there are gaps between the rates of growth of vectors tangent to the manifold and the rates of growth in the stable and unstable bundles that span the normal bundle. See [Fenichel71, Fenichel74, Fenichel77, Pesin, BatesLZ08], and also Appendix A.
@@
It should be noted that the unboundedness of manifolds is not merely a technical inconvenience; it can give rise to new geometric phenomena, such as an NHIM folding into itself. See [Eldering12, Example 3.8], reproduced here in Figure 2. To avoid such pathological examples, we have introduced an explicit assumption (U2) that the NHIM has a uniform tubular neighborhood.
@@ 1.2.4. Scattering maps for conformally symplectic systems
An important tool in the study of homoclinic excursions to a NHIM is the scattering map, introduced in [DLS00].
We consider a class of homoclinic excursions given by a transversal intersection of stable/unstable manifolds (the precise conditions are given in Definition 2.22 in Section 2.10). Given a homoclinic excursion $\{f^{n}(x)\}_{n\in\mathbb{Z}}$ to a NHIM $\Lambda$, there are unique points $x_{+},x_{-}\in\Lambda$ such that
@@
where $\lambda$ is strictly bigger than the rates in the tangent space to the NHIM.
@@
The scattering map is a very useful tool to understand long range motions and instability. Most of the applications of the scattering map in the instability problem have been for symplectic dynamics, since it gives a global way to connect invariant objects such as whiskered tori of (possibly) different topology or dimension [DelshamsLS06a, zbMATH06567884]. There are also results using the scattering map in systems with dissipation [Granados17, GideaLM22, AkingbadeGS23].
The scattering map is not itself part of the dynamics; rather, it should be regarded as a comparison between the dynamics restricted to the NHIM and the dynamics along the homoclinic excursion. Nevertheless, a surprising result [GideaLS20a, GideaLS20] shows that iterations of the scattering map interspersed with iterations of the dynamics restricted to the NHIM corresponds to true orbits of the map. The basic results of the previous two papers apply to general systems, including conformally symplectic systems. See also [DelshamsGR12].
Geometric properties of the scattering map in the symplectic case were systematically studied in [DLS08]. In this paper we develop similar results for conformally symplectic systems, but we encounter significant differences.
@@
    The formulas consist of a finite series plus an integral term as the remainder. Similar formulas were known for symplectic twist maps, and used for numerical computations [Tabacman95].
    The theory behind these formulas, in the case of unbounded manifolds, presents some notable differences from that of bounded manifolds, which are commonly considered in symplectic systems. As we pointed out, there are different action forms corresponding to gauge transformations. The gauges chosen can affect the convergence of the series.
@@
  • We provide a variational interpretation of the primitive function of the scattering map. We remark that, under certain convexity assumptions and provided the Legendre transform can be implemented, the primitive function of the scattering map coincides with the renormalized variational principle used to construct connecting orbits. [Rabinowitz08, Bessi96].
@@ 2.1.1. Riemannian manifolds, differentiable maps and forms
The sets of $\mathcal{C}^{r}$ vector fields, forms, and other objects with a linear structure, have a $\mathcal{C}^{r}$ norm given by the supremum of the size of the objects and their derivatives, which makes them into a Banach space. Under suitable uniformity assumptions (see (U1)), the set of $\mathcal{C}^{r}$ diffeomorphisms is a Banach manifold [Banyaga97].
@@ 2.1.2. Distances and regularity of manifolds
The following definitions are standard, but we record them here since there are subtle. See also [PughSW97].
@@
In this paper, we will need to study some foliations of a manifold by submanifolds. Foliations have regularity characterized by two numbers. One of them is the regularity of the leaves and another one is the dependence of the leaves on the base point along a transversal. We use the definition of [HPS, PughSW97].
@@ 2.1.3. Exponentials and connectors
It is well known [BurnsG05, Prop 4.5.2] that on a ball $B_{\rho_{x}}(0)\subseteq T_{x}M$ of radius $0<\rho_{x}\ll 1$, the exponential mapping is a diffeomorphism and induces a system of smooth coordinates on the neighborhood $\exp_{x}(B_{\rho_{x}}(0))$ of $x$ in $M$, referred to as geodesic coordinates. For any $x\in M$, the supremum of all such radii $\rho_{x}$ is called the injectivity radius.
@@
If $x,y$ are sufficiently close (depending only on derivatives of the metric) then $S_{x}^{y}$ is invertible, and we think of it as providing an identification of the two tangent spaces. The maps $S_{x}^{y}$ are called connectors in [HirschPPS69], where they are used in applications to hyperbolic systems.
@@ 2.2. A standing assumption on the manifold
Hypothesis (U1’) is sufficient but not necessary to obtain (U1). To obtain that the geodesic flow (and the exponential mapping) are $\mathcal{C}^{r}$, one does not need to assume that all the derivatives of $g$ are bounded. It suffices to assume only that the derivatives of the Riemann tensor are bounded. This property of a manifold is referred to as having bounded geometry [Cheeger1982finite, Eldering12].
@@ A $2$-form on a $\mathcal{C}^{r}$-manifold $M$ of even dimension $d$
The form $\omega$ is exact if $\omega=d\alpha$ for some $1$-form [^2] We refer to $\alpha$ as an action form for $\omega$ following [Haro00], but other names are used in the literature such as symplectic potential, Liouville form, or primitive form. $\alpha$ (not necessarily unique).
@@
A submanifold $L\subset M$ is said to be isotropic if $\omega_{\mid L}=0$, that is, for each $y\in L$, $T_{y}L\subseteq T_{y}L^{\omega}$, where $T_{y}L^{\omega}=\{v\in T_{y}M\,|\,\omega(y)(v,u)=0,\,\forall u\in T_{y}L\}$ is the symplectic orthogonal of $T_{y}L$. A submanifold $L\subset M$ is said to be coisotropic if for each $y\in L$, $T_{y}L^{\omega}\subseteq T_{y}L$. A submanifold $L\subset M$ is Lagrangian if it is both isotropic and coisotropic, or, equivalently, $\omega_{\mid L}=0$ and $\textrm{dim}(L)=\textrm{dim}(M)/2$. (See, e.g., [Weinstein71, Weinstein73].)
We will also pay attention to forms that are degenerate and have constant rank (dimension of the kernel) [Souriau97].
@@
In some works [Souriau97, LibermannM87], the definition of presymplectic forms also includes the constant rank condition that $\dim K_{x}(\omega)=$ const. on open sets. In general, the constant rank is not an open condition, since a small perturbation of the form may decrease the dimension of the kernel.
@@ A $\mathcal{C}^{r}$-map $f:M\to M$ is conformally symplectic if
It is well known([LibermannM87, Banyaga02]) that if $\omega$ is a two form of rank greater or equal than $4$, then, if, for a $1$-form $\beta$, we have $\beta\wedge\omega=0$, we conclude $\beta=0$.
@@
Applying $d$ to (2.3), $0=f^{*}(d\omega)=d(f^{*}\omega)=d\eta\wedge\omega$. Hence $d\eta=0$.
When $\eta=1$ a system satisfying (2.3) is symplectic, so the results in this paper (that do not include explicitly $\eta\neq 1$) imply the results in [DLS08]. When $\eta\neq 1$, the symplectic volume contracts or expands, and, therefore, invariant manifolds have volume zero or infinite. The interesting case is when the symplectic volume is infinite. In many mechanical applications, the physical friction satisfies $\eta<1$, so we will take this as the default case.
@@ 2.5. Exact conformally symplectic maps
In many cases of interest, there are topological reasons why the symplectic form $\omega$ has to be exact; see [ArnaudF24] and Section 4.1.
@@ 2.5.1. Gauge transformations
In general, if a map is exact for $\alpha$, it will also be exact under the gauge transformation of the action form  [^3] In electromagnetism, where the electromagnetic field is an exact 2-form exterior derivative of an electromagnetic 1-form (called vector potential), one often uses adding gradient to the vector potential [Thirring, Zangwill]. The term “gauge” was introduced in [Weyl51] for electromagnetic vector potentials
@@ 2.7. Normally hyperbolic invariant manifolds
Definition 2.16 says that in a small neighborhood of $\Lambda\subset M$ of any point $x\in\Lambda$, the manifold $M$ is the product of the manifold $\Lambda$ and of the fibers of the bundles $E^{s}$ and $E^{u}$. For compact manifolds, this gives a tubular neighborhood of $\Lambda$. See [HPS].
For unbounded manifolds, it can happen that the manifold $\Lambda$ folds back and comes close to itself. A concrete example appears in [Eldering12, Example 3.8], which we illustrate in Fig. 2. The content of assumption (U2) is that this folding back of the manifold $\Lambda$ does not happen.
Figure 2. A NHIM without a uniform tubular neighborhood. After [Eldering12, Example 3.8]
@@
(2.15) $W^{u,s,\textrm{loc}}_{\Lambda}=\bigcup_{x\in\Lambda}W^{s,u,\textrm{loc}}_{x}.$
The union on the right-hand side of (2.15) is a disjoint union. Therefore, the decomposition of $W^{u,s,\textrm{loc}}_{\Lambda}$ into strong stable leaves is a foliation and, moreover, the leaves are identified uniquely by their intersection with $\Lambda$.
These (un)stable and strong (un)stable manifolds and their regularity are described in Appendix A (see Theorem A.1). We will take advantage of the remarkable fact that $W^{u,s,\textrm{loc}}_{\Lambda}$ admit topological characterizations, but has consequences for rates of convergence and regularity.
@@
The assumption (N) is to avoid complicated statements. If $\mu_{+}<1$, there is a unique fixed point $p\in\Lambda$, and $\Lambda$ is a submanifold of the stable manifold of $p$. Many of the regularity statements obtained in the general theory of NHIMs remain true, but they are far from optimal. Also, in some estimates (derived from [DLS08, Proposition 15] or similar) we use bounds like $C_{1}\mu_{+}^{n}+C_{2}\mu_{+}^{2n}\leq C\mu_{+}^{2n}$, for $n\geq 0$. If we do not assume (N), different algebraic expressions for the bounds would be required for $\mu_{+}\geq 1$ and for $\mu_{+}<1$, and therefore many statements would need to distinguish between the two different cases.